Bridgeport Corp. had 1,800 units of part T on hand April 1, 2017, costing $8.00 each. During April, Bridgeport made the following purchases of part T.

Units Unit
Cost
April 4 2,900 $8.25
10 5,000 8.40
19 1,200 9.00
29 2,800 9.90

A physical count at April 30, 2017 showed 4,300 units of Part T on hand.

Using the FIFO method, what is the cost of part T inventory at April 30, 2017? Using the LIFO method, what is the inventory cost? Using the average-cost method, what is the inventory cost?
Answer #1
Unit Unit Cost Total Cost
Beginning inventory 1800 8 14400
April 4 2900 8.25 23925
April 10 5000 8.40 42000
April 19 1200 9 10800
April 29 2800 9.90 27720
Total 13700 118845

Calculate Ending inventory :

FIFO = 27720+10800+(300*8.4) = 41040

LIFO = 14400+(2500*8.25) = 35025

Average cost = 118845/13700*4300 = $37302